OpenStudy (anonymous):

IM GIVING MEDALS FOR THE BEST ANSWER The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2)

OpenStudy (anonymous):

y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

I mean like to the person that hep me understand it the best

OpenStudy (anonymous):

what do you think parallel is?

OpenStudy (anonymous):

i know what parallel is

OpenStudy (anonymous):

ok, so it has same slope, correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes it has a slope

OpenStudy (cydney_morgan):

Both guys here deserve a medal... but honestly if you help people JUT=ST so you can get medals thats pretty sad, you shoud be helping people out of the goodness of your heart <3 I'll give the medal to someone, and I hope the owner of this question will give the medal to the other person.

OpenStudy (anonymous):

so which one has the same slope? it should not have the exact same equation.

OpenStudy (anonymous):

@Cydney_Morgan u are right :D

OpenStudy (anonymous):

well I like to help with my yes ok then lets get started

OpenStudy (anonymous):

yes she is

OpenStudy (anonymous):

@R00nnYBraiNsbiG what do you think?

OpenStudy (anonymous):

about your question...

OpenStudy (anonymous):

let me see doing reseach

OpenStudy (anonymous):

tbh i still dont know which answer it is

OpenStudy (cydney_morgan):

@MelissaHolmes Ty <3

OpenStudy (anonymous):

:D yw

OpenStudy (anonymous):

Write the equation for the line in slope-intercept form: Slope: m=4/3; Y-intercept: b =-2 y = 4/3x - 2 y = 1/2x - 2 ... Find the line that passes through the point (0, 2) and is parallel to the line y = -7/3x - 4. y ... Choose the graph that represents a line that contains the point (-3,2) and has a slope of 1/2. a b c 21.

OpenStudy (anonymous):

and so which two have a slope of 2?

OpenStudy (anonymous):

ok well thanks guys yall are the best

OpenStudy (anonymous):

. Given the line 2x - 3y = 9 and the point (4, -1), find lines through the point that are ... 2x - 3y = 9 -3y = -2x + 9 y = (2/3)x - 3 seems correct?

OpenStudy (anonymous):

ok... so u figure out the answer?

OpenStudy (anonymous):

and no @TICOxTICO it has to have the same slope to be parallel

OpenStudy (anonymous):

Okay seems you deserve the best response that you have bested me for that I respect that.

OpenStudy (anonymous):

ok... but you gave it your best, so I appreciate that also :D

OpenStudy (anonymous):

yes i did thanks

OpenStudy (anonymous):

Yes it is most important to try =D

OpenStudy (anonymous):

ok @R00nnYBraiNsbiG that's great! :D

OpenStudy (anonymous):

you are most welcome from me and MelissaHolmes